# Bound States and Particle Production by Breather-Type Background Field Configurations

### v1

@@ Abstract
We investigate the interaction of fermion fields with oscillating domain walls, inspired by breather-type solutions of the sine-Gordon equation. Our study focuses on the fermionic bound states and particle production induced by a time-dependent scalar background field. The fermions couple to two domain walls undergoing harmonic motion, and we explore the resulting dynamics of the fermionic wave functions. We demonstrate that while fermions initially form bound states around the domain walls, the energy provided by the oscillatory motion of the scalar field induces an outward flux of fermions and antifermions, leading to particle production and eventual flux propagation toward spatial infinity. Through numerical simulations, we observe that the fermion density exhibits quasiperiodic behavior, with partial recurrences of the bound state configurations after each oscillation period. However, the fermion wave functions do not remain localized, and over time, the density decreases as more particles escape the vicinity of the domain walls. Our results highlight that the sine-Gordon-like breather background, when coupled to fermions, does not preserve integrability or stability, with the oscillations driving a continuous energy transfer into the fermionic modes. This study sheds light on the challenges of maintaining steady-state fermion solutions in time-dependent topological backgrounds and offers insights into particle production mechanisms in oscillating soliton systems.
@@ 1 Introduction
Particle production by extended classical structures is a topic that has been studied since the earliest days of quantum mechanics. In nontrivial spacetime backgrounds, this has been an incredibly fruitful line of inquiry, with the theoretical discoveries of the Unruh effect (for an accelerating observer) [1] and Hawking radiation (in a black hole spacetime) [2] having fundamentally changed our understanding of spacetime physics. The creation of particles can be seen as a process of transitions between different Fock states of a quantum theory embedded in an interesting classical background, and the relevant background can involve fields other than the spacetime metric. In particular, the coupling of particles to classical scalar and vector fields can have many intriguing consequences.
@@
When the background involves an extended structure built out of a bosonic field, the coupling of one or more additional fermion fields can be particular interesting. Even without particle creation, the coupling of fermions to a topologically nontrivial bosonic field configuration can lead to the fascinating phenomenon of fermion fractionalization [3]. The nontrivial topology of the background may produce a potential for the fermion field whose spectrum is guaranteed to include a nondegenerate zero mode. When the fermion field is quantized, with the Dirac sea of negative-energy states filled, the zero mode is assigned partially to the Dirac sea and partially to the positive-energy part of the spectrum. The level in the energy spectrum at with the fermion number $n_{F}$ vanishes actually lies in the middle of the zero mode. Since the zero mode cannot ever actually be only partially filled, the real Fock states (with the mode filled or unfilled) consequently have non-integer values of $n_{F}$.
@@
The situation may be even more complicated when there are multiple solitary wave structures present in the background. In the presence of multiple stationary solitons, there are typically multiple zero-energy (or almost-zero-energy) modes as part of the fermionic spectrum. When the solitary wave features are in motion relative to one-another, this is further augmented by the possibility of fermion or antifermion particle creation, and it is a type of scenario like this that will be the focus of this paper. This is part of an ongoing research program [4], which has so far examined the fermion spectrum in backgrounds with two or three solitary waves. However, one of the challenges to doing calculations in a theory like this—in which the energy differences between fermion states bound to bosonic solitary wave may be exponentially small—is that the background may not be known precisely. Solitary waves are a feature of nonlinearly coupled classical fields, and in most cases exact solutions involving multiple topological excitations are either unknown or extraordinarily complicated and impossible to express concisely. Small discrepancies between the exact energies of the background field configurations and the approximate forms used for calculational expediency may become a problem when trying to study the very small energy differences between neighboring fermion states.
@@
In this paper, we shall look at fermion solutions in a particular type of background, inspired by the sine-Gordon equation in $1+1$ dimensions, since the sine-Gordon equation has time-dependent, multiple-solitons solutions that are known exactly. The focus will continue to be on the manifold of almost-zero-energy states that generalize the Jackiw-Rebbi zero modes. In section 2, we shall review the sine-Gordon equation and its breather-type soliton solutions. The consideration of a breather background was motivated by the hope that the breathers, which are bound soliton-antisoliton states with periodic time dependence, might support fermion bound states that were similarly periodic and for which exact solutions might be even be found. However, solving the Dirac equation in this kind of time-dependent background still turns out to be too difficult to do analytically, so in section 3, we introduce another kind of bosonic background, inspired by the sine-Gordon breather but involving step functions that make the fermionic theory amenable to additional analytic solutions methods. In section 4 we show, using a Fourier decomposition of the fermionic modes, that for neither the step function background, nor for the true sine-Gordon breather background, can there actually be steady-state fermion solutions. There is an unavoidable outward flux in the fermion and antifermion nodes, driven energetically by the oscillations of the background scalar field. In section 5, we discuss the computational challenges this poses and present numerical results for how fermionic states that start off closely bound to the solitary waves in the background produce wave of flux escaping to spatial infinity. Section 6 gives our final conclusions.
@@ 2 Breather Solitons
The theory this describes is extremely special, having degenerate vacua at at all $\phi=n$ and solitary wave solutions that interpolate between them. These solitary waves are quite unusual, since they can pass through one-another with no net deformations or radiation once they are well separated again. The solitary waves are thus actually “solitons” in the oldest and strongest sense of that term. The theory is integrable, a particularly unusual property in a relativistic theory and one that is closely connected to the solitary wave structure. The schematic (but nonrigorous) argument that the presence of true solitons means that there must be an infinite number of conservation laws is fairly straightforward Suppose that $N$ solitons are moving along $x$-axis; any $N\geq 0$ is possible. The can all have different speeds, so that as $t\rightarrow\infty$ the separations between the solitary waves will become arbitrarily large, and thus the fraction of the total energy that remains tied up in the interactions between the solitons shrinks arbitrarily small. Each soliton is essentially an isolated system with conserved internal and kinetic energies of its own; and since $N$ may be arbitrarily large, there must likewise be an unlimited number of these conservation laws. Such localized single-soliton energies are quite different from the usual way in which the infinite family of sine-Gordon conservation laws are normally expressed. However, the motivation for this argument was simply to show the deep connection between the existence of many-soliton solutions for $\phi$ and existence of an infinite number of conservation laws.
A very natural question to ask is how many of the special properties of the sine-Gordon theory persist when the bosonic theory is coupled to a Dirac fermion field. Solutions with single solitary waves coupled to the fermions are well understood, with the unusual observation being the presence of zero-energy bound state modes and the surprising consequence that the fermion number actually has fractional eigenvalues [3]—among other anomalous features [5, 6, 7, 8]. Studying fermions in scalar backgrounds composed of multiple solitary wave is a complicated subject, however. Normally, it is only amenable to approximate analyses—whether analytical or numerical [4, 9, 10, 11, 12]. However, one might hope that the simplicity of the sine-Gordon system would make it more straightforward to locate exact solutions of the coupled boson-fermion system. Unfortunately, this does not seem to be the case, and, in fact, the first step in our analysis will be to replace the exact sine-Gordon background with a seemingly artificial approximation. Nonetheless, we shall find that our results can reveal quite a bit about the general structure of the Dirac theory in a time-dependent oscillatory topological background.
@@
This is the “natural” form for the coupling of the Dirac field to the scalar, because with a properly chosen value of $g$, the action becomes supersymmetric. In that case, the interactions in the boson and fermion Lagrange densities each descend from a “superpotential” $W(\phi)=\sin(\pi\phi/2)$, since $(1-\cos\pi\phi)\propto W^{2}$ and $\cos(\pi\phi/2)\propto dW/d\phi$. (However, Ref. [11] used an ordinary Yukawa coupling even when the bosons obeyed a sine-Gordon equation.)
@@ 3 Dirac Equation in the Kink-Antikink Background
Unfortunately, no exact solutions to the Dirac equation in the breather background have been forthcoming. So instead of the exact sine-Gordon breather solution, we shall take two step function domain walls undergoing simple harmonic motion back and forth through one-another, with amplitude $v\geq 0$ and frequency $\omega$,
@@
In the presence of a well-separate kink-antikink pair, there are two almost-zero-energy fermion modes, with even and odd behaviors under parity. When the domain walls are far apart, the corresponding fermion wave function are well approximated by symmetric and antisymmetric linear combinations of a purely-$\psi_{1}$ bound state wave function tightly localized around one domain wall’s location and a purely-$\psi_{2}$ function localized around the other solitary wave. In this regime, the two fermion states are approximately equally displaced above and below zero energy. In contrast, when the kink and antikink draw close together, so that $v|\cos\omega t|\lesssim g^{-1}$, the wave functions become much more complicated. Both $\psi_{1}$ and $\psi_{2}$ are appreciable in the neighborhood of $x=0$, and their energy shifts are no longer equal and opposite. Moreover, even though the kink and antikink pass through one-another, the almost-zero-energy modes do not follow along. After their episodes of complicated mixing during each interval in which $v|\cos\omega t|\lesssim g^{-1}$, the $\psi_{1}$-dominated and $\psi_{2}$-dominated portions of the full wave function do not pass through each other but instead rebound, switching their locations of attachments from kink to antikink and vice versa. The wave function $\psi$ is periodic with period $\pi/\omega$, like $V$, rather than the $2\pi/\omega$ periodicity of $\phi$ itself. Moreover, the evolution of the system during the collision period cannot be adiabatic; although the adiabatic approximation may be excellent when the kink and antikink are widely separated, the field backgrounds (2) or (4) both pass through the state $\phi=0$—which is a background in which the discrete modes of the instantaneous Hamiltonian merge into a massless continuum that extends through $E=0$ with no gap.
@@ 4 Fourier Modes of the Dirac Equation
The observation that a steady-state, $(\pi/\omega)$-periodic solution of the Dirac-equation necessarily involves a outgoing flux of quanta from the kink-antikink region suggests that the one of the original reasons for focusing especially on the sine-Gordon breather system was misaimed. The purely bosonic sine-Gordon theory is integrable, with an infinite number of conserved quantities. Recall, the solitary waves are true “solitons” in the strong sense, in that while they exert net forces on one-another, when the collide they will pass through one-another and as they separate to great distances, return to their original shapes. In particular, while there are the breather solutions, in which a kink and an antikink form a bound state, never will a kink and an antikink annihilate into non-topological low-amplitude radiation excitations. It was hoped that including appropriately coupled fermions in the theory (perhaps in the supersymmetric version of the boson-fermion theory) might preserve the integrability of the theory and the absolute stability of individual solitary waves. However, the argument for the structure (24) for the large-$n$ Fourier modes’ behavior in the far field does not depend in any fundamental way on the specific step function approximation for the $\phi$ background; the reasoning would apply equally well with (2) as with (4). So the fact that the solutions of the Dirac equation in a breather-like background involve outward net flux suggests that any hope that the sine-Gordon theory coupled to fermions might remain fully integrable appears to be untenable.
@@ 5 Computational Complications and Results
The fact that fermionic modes can propagate outside the kink-antikink region (completely freely in the square wave model, nearly freely in the more realistic sine-Gordon theory) poses numerical as well as analytical challenges. An algorithm for integrating the Dirac equation is going to encounter difficulties at the boundaries of the integration domain, because the solutions of the unmodified Dirac equation may be dominated by convective behavior—which will typically involve energies modes propagating inward from infinity as well as outward toward infinity. This is obviously contrary to the radiation boundary conditions we would like to impose.
If the initial Dirac wave function is well localized around the soliton-containing region, then at small times the numerical solution in the radiation zone will be dominated, as desired, by the outgoing fermion and antifermion modes. However, if the Dirac field is not subject to unphysical conditions at or near the ends of the integration domain, then when the outgoing excitations start to impinge on the exterior boundaries, the algorithm will suffer from an unstable explosion of unphysical incoming mode excitations. The field itself will begin to blow up at the boundary, and the inward propagation will rapidly wreck the calculated solution.
@@
Although initially $\psi_{1}$ is strongly localized around the kink and $\psi_{2}$ around the antikink, the time-dependence of the background causes the solution around $x_{K}=-v=-10$ to acquire a $\psi_{2}$ component fairly quickly. The kink and antikink in (4) start from rest, but it does not take long for their motion to couple the two wave function components together. As a result, over most of the integration time interval, $\psi_{1}$ and $\psi_{2}$ are comparable everywhere in the spatial domain. From each of the regions around $x=\pm v$ where the wave functions are initially localized, each of $\psi_{1}$ and $\psi_{2}$ shows an pattern of expansion at the speed $c_{s}$. After a time $v/c_{s}=10$, the expanding cones overlap, and there is an interference region centered on $x=0$. Interestingly, as a consequence of the way the equations (1011) are coupled (with the coupling terms having explicit factors of $i$), the contribution to $\psi_{2}$ in the initial vicinity of $\psi_{1}$ (and vice versa) is mostly imaginary, while $\psi_{1}$ in that region remains predominantly real—at least until the expanding light cones based at ±$\pm v$ begin to overlap. This fact is evident from figure 4, which shows the typical behavior of the real and imaginary parts of $\psi$.

### v2

@@ Abstract
We investigate the interaction of fermion fields with oscillating domain walls, inspired by breather-type solutions of the sine-Gordon equation, a nonlinear system of fundamental importance. Our study focuses on the fermionic bound states and particle production induced by a time-dependent scalar background field. The fermions couple to two domain walls undergoing harmonic motion, and we explore the resulting dynamics of the fermionic wave functions. We demonstrate that while fermions initially form bound states around the domain walls, the energy provided by the oscillatory motion of the scalar field induces an outward flux of fermions and antifermions, leading to particle production and eventual flux propagation toward spatial infinity. Through numerical simulations, we observe that the fermion density exhibits quasiperiodic behavior, with partial recurrences of the bound state configurations after each oscillation period. However, the fermion wave functions do not remain localized, and over time, the density decreases as more particles escape the vicinity of the domain walls. Our results highlight that the sine-Gordon-like breather background, when coupled non-supersymmetrically to fermions, does not preserve integrability or stability, with the oscillations driving a continuous energy transfer into the fermionic modes. This study sheds light on the challenges of maintaining steady-state fermion solutions in time-dependent topological backgrounds and offers insights into particle production mechanisms in nonlinear dynamical systems with oscillating solitons.
@@ 1 Introduction
Particle production by extended classical structures is a topic that has been studied since the earliest days of quantum mechanics. In nontrivial spacetime backgrounds, this has been an incredibly fruitful line of inquiry, with the theoretical discoveries of the Unruh effect (for an accelerating observer) [1] and Hawking radiation (in a black hole spacetime) [2] having fundamentally changed our understanding of spacetime physics. The creation of particles can be seen as a process of transitions between different Fock states of a quantum theory embedded in an interesting classical background, and the relevant background can involve fields other than the spacetime metric. In particular, the coupling of quantum-mechanical particles particles to classical scalar and vector fields can have many intriguing consequences [3, 4, 5, 6, 7, 8, 9]. The nonlinear interactions responsible for both the shapes of solitary waves and the effective couplings between multiple solitary waves [10, 11, 12, 13, 14, 15] are significant topics of research by themselves. For instance, during a sufficiently energetic collision of solitary wave pair in $1+1$ dimensions, there may be a strongly inelastic bounce—in which one or both of the solitary waves rebound back from the impact with internal modes or continuum scalar field modes excited. Over long times, the behavior of repeatedly colliding and rebounding solitary waves may display an intricate “fractal” dependence on the initial conditions [16, 17, 18, 19, 20, 21].
@@
However, when the background involves an extended structure built out of a bosonic field, the coupling of one or more additional fermion fields can be considered particularly interesting. Even without particle creation, the coupling of fermions to a topologically nontrivial bosonic field configuration can lead to the fascinating phenomenon of fermion fractionalization [3]. The nontrivial topology of the background may produce a potential for the fermion field whose spectrum is guaranteed to include a nondegenerate zero mode. When the fermion field is quantized, with the Dirac sea of negative-energy states filled, the zero mode is assigned partially to the Dirac sea and partially to the positive-energy part of the spectrum. The level in the energy spectrum at which the fermion number $n_{F}$ vanishes actually lies in the middle of the zero mode [22]. Since the zero mode cannot ever actually be only partially filled, the real Fock states (with the mode filled or unfilled) consequently have non-integer values of $n_{F}$ [23, 24].
@@
The situation may be even more complicated when there are multiple solitary wave structures present in the background. In the presence of multiple stationary solitons, there are typically multiple zero-energy (or almost-zero-energy) modes as part of the fermionic spectrum. When the solitary wave features are in motion relative to one-another, this is further augmented by the possibility of fermion or antifermion particle creation [25], and it is a type of scenario like this that will be the focus of this paper. This is part of an ongoing research program [26], which has so far examined the fermion spectrum in backgrounds with two or three solitary waves. However, one of the challenges to doing calculations in a theory like this—in which the energy differences between fermion states bound to bosonic solitary wave may be exponentially small—is that the background may not be known precisely. Solitary waves are a feature of nonlinearly coupled classical fields, and in most cases analytic solutions involving multiple topological excitations are either unknown or extraordinarily complicated and impossible to express concisely. Small discrepancies between the precise energies of the background field configurations and the approximate forms used for calculational expediency may become a problem when trying to study the very small energy differences between neighboring fermion states.
@@
In this paper, we shall look at fermion solutions in a particular type of background, inspired by the sine-Gordon equation in $1+1$ dimensions, since the sine-Gordon equation has time-dependent, multiple-solitons solutions that are known analytically. The focus will continue to be on the manifold of almost-zero-energy states that generalize the Jackiw-Rebbi zero modes; our work represents new fundamental advances in this area of investigation. In section 2, we shall review the sine-Gordon equation and its breather-type soliton solutions. The consideration of a breather background was motivated by the hope that the breathers, which are bound soliton-antisoliton states with periodic time dependence, might support fermion bound states that were similarly periodic and for which analytic solutions might be even be found. However, solving the Dirac equation in this kind of time-dependent background still turns out to be too difficult to do analytically, so in section 3, we introduce another kind of bosonic background, inspired by the sine-Gordon breather but involving step functions that make the fermionic theory amenable to additional analytic solutions methods. In section 4 we show, using a Fourier decomposition of the fermionic modes, that for neither the step function background, nor for the true sine-Gordon breather background, can there actually be steady-state fermion solutions. There is an unavoidable outward flux in the fermion and antifermion nodes, driven energetically by the oscillations of the background scalar field. This is an entirely new result. In section 5, we discuss the computational challenges this poses and present numerical results for how fermionic states that start off closely bound to the solitary waves in the background produce wave of flux escaping to spatial infinity. Section 6 gives our final conclusions.
@@ 2 Breather Solitons
The theory this describes is extremely special, having degenerate vacua at all $\phi=n$ and solitary wave solutions that interpolate between them. These solitary waves are quite unusual, since they can pass through one-another with no net deformations or radiation once they are well separated again. The solitary waves are thus actually “solitons” in the oldest and strongest sense of that term. The theory is integrable, a particularly unusual property in a relativistic theory and one that is closely connected to the solitary wave structure. The schematic (but nonrigorous) argument that the presence of true solitons means that there must be an infinite number of conservation laws is fairly straightforward Suppose that $N$ solitons are moving along $x$-axis; any $N\geq 0$ is possible. They can all have different speeds, so that as $t\rightarrow\infty$ the separations between the solitary waves will become arbitrarily large, and thus the fraction of the total energy that remains tied up in the interactions between the solitons shrinks arbitrarily small. Each soliton is essentially an isolated system with conserved internal and kinetic energies of its own; and since $N$ may be arbitrarily large, there must likewise be an unlimited number of these conservation laws. Such localized single-soliton energies are quite different from the usual way in which the infinite family of sine-Gordon conservation laws are normally expressed. However, the motivation for this argument was simply to show the deep connection between the existence of many-soliton solutions for $\phi$ and existence of an infinite number of conservation laws. (Related analyses of integrable equations in $2+1$ dimensions are also currently an active area of study [27, 28, 29, 30].)
A very natural question to ask is how many of the special properties of the sine-Gordon theory persist when the bosonic theory is coupled to a Dirac fermion field. Solutions with single solitary waves coupled to the fermions are well understood, with the unusual observation being the presence of zero-energy bound state modes and the surprising consequence that the fermion number actually has fractional eigenvalues [3]—among other anomalous features [31, 32, 33, 34]. Studying fermions in scalar backgrounds composed of multiple solitary wave is a complicated subject, however. Normally, it is only amenable to approximate analyses—whether analytical or numerical [26, 35, 36, 37, 38]. However, one might hope that the simplicity of the sine-Gordon system would make it more straightforward to locate complete analytic solutions of the coupled boson-fermion system. Unfortunately, this does not seem to be the case, and, in fact, the first step in our analysis will be to replace the sine-Gordon breather background with a seemingly artificial approximation. Nonetheless, we shall find that our results can reveal quite a bit about the general structure of the Dirac theory in a time-dependent oscillatory topological background.
@@
This is the “natural” form for the coupling of the Dirac field to the scalar, because with a properly chosen value of $g$, the action becomes supersymmetric. In that case, the interactions in the boson and fermion Lagrange densities each descend from a “superpotential” $W(\phi)=\sin(\pi\phi/2)$, since $(1-\cos\pi\phi)\propto W^{2}$ and $\cos(\pi\phi/2)\propto dW/d\phi$. (However, Ref. [37] used an ordinary Yukawa coupling even when the bosons obeyed a sine-Gordon equation.) Perhaps ironically, however, we do not actually expect our general conclusions to apply in the precisely supersymmetric sine-Gordon model, which remains integrable—like the purely bosonic model [39, 40]. The existence of an infinite number of degrees of freedom can qualitatively change the behavior of the system.
@@ 3 Dirac Equation in the Kink-Antikink Background
Unfortunately, no analytic solutions to the Dirac equation in the breather background have been forthcoming. So instead of the anharmonically-oscillating sine-Gordon breather solution, we shall take two step function domain walls undergoing simple harmonic motion back and forth through one-another, with amplitude $v\geq 0$ and frequency $\omega$,
@@
In the presence of a well-separated kink-antikink pair, there are two almost-zero-energy fermion modes, with even and odd behaviors under parity. When the domain walls are far apart, the corresponding fermion wave function are well approximated by symmetric and antisymmetric linear combinations of a purely-$\psi_{1}$ bound state wave function tightly localized around one domain wall’s location and a purely-$\psi_{2}$ function localized around the other solitary wave. In this regime, the two fermion states are approximately equally displaced above and below zero energy. In contrast, when the kink and antikink draw close together, so that $v|\cos\omega t|\lesssim g^{-1}$, the wave functions become much more complicated. Both $\psi_{1}$ and $\psi_{2}$ are appreciable in the neighborhood of $x=0$, and their energy shifts are no longer equal and opposite. Moreover, even though the kink and antikink pass through one-another, the almost-zero-energy modes do not follow along. After their episodes of complicated mixing during each interval in which $v|\cos\omega t|\lesssim g^{-1}$, the $\psi_{1}$-dominated and $\psi_{2}$-dominated portions of the full wave function do not pass through each other but instead rebound, switching their locations of attachments from kink to antikink and vice versa. The wave function $\psi$ is periodic with period $\pi/\omega$, like $V$, rather than the $2\pi/\omega$ periodicity of $\phi$ itself. Moreover, the evolution of the system during the collision period cannot be adiabatic; although the adiabatic approximation may be excellent when the kink and antikink are widely separated, the field backgrounds (2) or (4) both pass through the state $\phi=0$—which is a background in which the discrete modes of the instantaneous Hamiltonian merge into a massless continuum that extends through $E=0$ with no gap.
@@ 4 Fourier Modes of the Dirac Equation
The observation that a steady-state, $(\pi/\omega)$-periodic solution of the Dirac-equation necessarily involves a outgoing flux of quanta from the kink-antikink region suggests that the one of the original reasons for focusing especially on the sine-Gordon breather system was misaimed. The purely bosonic sine-Gordon theory is integrable, with an infinite number of conserved quantities. Recall, the solitary waves are true “solitons” in the strong sense, in that while they exert net forces on one-another, when the collide they will pass through one-another and as they separate to great distances, return to their original shapes. In particular, while there are the breather solutions, in which a kink and an antikink form a bound state, never will a kink and an antikink annihilate into non-topological low-amplitude radiation excitations. It was hoped that including appropriately coupled fermions in the theory (perhaps in the version of the boson-fermion theory which differs from the supersymmetric theory only in the magnitude of the coupling constant) might preserve the integrability of the theory and the absolute stability of individual solitary waves. However, the argument for the structure (24) for the large-$n$ Fourier modes’ behavior in the far field does not depend in any fundamental way on the specific step function approximation for the $\phi$ background; the reasoning would apply equally well with (2) as with (4). Any one of the large-$n$ modes should generally be excited, unless there is a conservation law that prohibits such an excitation. This is why the choice of coupling that makes the fermion-coupled sine-Gordon model truly supersymmetric differs from all other values. The supersymmetric version has an infinite number of conservation laws and is thus capable of preventing the excitation of the infinite number of outward-propagating modes and preserving perpetual breathers. However, the fact that the solutions of the Dirac equation in a breather-like background otherwise involve outward net flux suggests that any hope that the most general sine-Gordon theory coupled to fermions might remain fully integrable appears to be untenable.
@@ 5 Computational Complications and Results
The fact that fermionic modes can propagate outside the kink-antikink region (completely freely in the square wave model, nearly freely in the more realistic sine-Gordon theory) poses numerical as well as analytical challenges. Because of the greater complexity of the sine-Gordon breather system, we shall continue to focus, as we move to numerical calculations, on the model with step function solitons. We shall also neglect the back-reaction of the fermions on the boson field, treating $\phi$ as a unmodified oscillating background. This is a natural approximation when the coupling $g$ is small, and is in any case necessary when working with the signum function domain walls, whose dynamical equations would involve singular quantities.
An algorithm for integrating the Dirac equation is going to encounter difficulties at the boundaries of the integration domain, because the solutions of the unmodified Dirac equation may be dominated by convective behavior—which will typically involve energies modes propagating inward from infinity as well as outward toward infinity. This is obviously contrary to the radiation boundary conditions we would like to impose.
If the initial Dirac wave function is well localized around the soliton-containing region, then at small times the numerical solution in the radiation zone will be dominated, as desired, by the outgoing fermion and antifermion modes. However, if the Dirac field is not subject to unphysical conditions (such as the stong spatial damping described below) at or near the ends of the integration domain, then when the outgoing excitations start to impinge on the exterior boundaries, the algorithm will suffer from an unstable explosion of unphysical incoming mode excitations. A small numerical error will introduce an admixture of incoming waves into the solution, with the error due to these spurious modes increasing with time. The field itself will begin to blow up at the boundary, and the inward propagation will rapidly wreck the calculated solution.
@@
Although initially $\psi_{1}$ is strongly localized around the kink and $\psi_{2}$ around the antikink, the time-dependence of the background causes the solution around $x_{K}=-v=-10$ to acquire a $\psi_{2}$ component fairly quickly. The kink and antikink in (4) start from rest, but it does not take long for their motion to couple the two wave function components together. As a result, over most of the integration time interval, $\psi_{1}$ and $\psi_{2}$ are comparable everywhere in the spatial domain. From each of the regions around $x=\pm v$ where the wave functions are initially localized, each of $\psi_{1}$ and $\psi_{2}$ shows an pattern of expansion at the speed $c_{s}$. At short times, this may look like oscillation-stimulated tunneling from the bound state to continuum states. However, the outflow from the soliton regions continues indefinitely; in particular, there continues to be outgoing radiation after the oscillation amplitude of $|\psi|^{2}$ in the immediate vicinity of the domain walls has fallen to be no larger than the typical values elsewhere inside the light cone. After a time $v/c_{s}=10$, the expanding cones overlap, and there is an interference region centered on $x=0$. Interestingly, as a consequence of the way the equations (1011) are coupled (with the coupling terms having explicit factors of $i$), the contribution to $\psi_{2}$ in the initial vicinity of $\psi_{1}$ (and vice versa) is mostly imaginary, while $\psi_{1}$ in that region remains predominantly real—at least until the expanding light cones based at ±$\pm v$ begin to overlap. This fact is evident from figure 4, which shows the typical behavior of the real and imaginary parts of $\psi$.